stat 16

Eight pairs of data yield r = 0.708 and the regression equation ^y= 55.8 + 2.79x.also _y=71.125.What is the best predicted value of y for x = 6.8 ?

a. 71.13

b. 74.77

c. 57.80

d. 382.23
Answer #1
Y = 55.8 + 2.79x.

for x = 6.8, y = 55.8 +(2.79*6.8) = 74.77

hence the correct answeris b. 74.77